What is the relationship between molefraction of solute(XB), molarity(M), density of solution(Ρ), mass of


solute (MB) &  mass of solvent (MA) . How to derive it?

Grade:12

1 Answers

Vikas TU
3 years ago

Molarity =moles of solute/litre of solution

Molality=moles of solute/kg of solute

Mole fraction=Number of moles of solute /Total moles of solution

Relation b/w molarity,molality and mole fraction

1/m = (d/M)-(Mb/1000)=1/1000[XaMa/Xb]

Where m=molality of solution

d= density of solution

M=molarity of solution

Mb=molar mass of solute

Xa=mole fraction of solvent

Xb=mole fraction of solute

Ma=molar mass of solvent